# Less Is More [Uncaptioned image]: Scalable Visual Navigation from Limited Data

### v1

@@ IV-D Ablations
Figure 5: (Left) Evaluation of SPL on $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{AUG}}$ test sets trained on a varying number of augmenting geometric paths. Adding geometric samples mainly benefits performance on the more diverse $\mathcal{D}_{\text{AUG}}$ test split. (Middle) When evaluating goals following the $\mathcal{D}_{\text{AUG}}$ distribution, models trained solely on $\mathcal{D}_{\text{TEL}}$ consistently underperform compared to those trained on the $\mathcal{D}_{\text{AUG}}$. While training on $\mathcal{D}_{\text{AUG}}$ generalizes well to $\mathcal{D}_{\text{TEL}}$-style goals, models trained only on $\mathcal{D}_{\text{TEL}}$ fail to generalize to unseen diverse goals with limited data. (Right) Evaluation of SPL on $\mathcal{D}_{\text{TEL}}$ goals trained using the $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{AUG}}$. Training on just 5 missions of $\mathcal{D}_{\text{AUG}}$ achieves comparable performance to training on $\mathcal{D}_{\text{TEL}}$ with 41 missions.
@@ IV-D1 Training with fewer geometric paths per frame
To explore the benefits of adding geometric paths, we train models on different mixtures of $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{GEO}}$. The policies are evaluated on the held-out test missions with $\mathcal{D}_{\text{TEL}}$ or more diverse $\mathcal{D}_{\text{AUG}}$ distributed goals. See Fig. 5 (left).
@@ IV-D2 Training on fewer missions
To understand how well policies generalize when trained with fewer missions, we compare performance across varying numbers of training missions drawn from either $\mathcal{D}_{\text{TEL}}$ or $\mathcal{D}_{\text{AUG}}$. Policies trained on a small number of diverse $\mathcal{D}_{\text{AUG}}$ missions retain strong performance on $\mathcal{D}_{\text{TEL}}$ goals, whereas policies trained only on $\mathcal{D}_{\text{TEL}}$ require substantially more missions to reach comparable SPL (see Fig. 5 (Right)). This highlights that diversity in the training data, rather than sheer quantity, is key for efficient generalization (Q1, Q2). Tab III in the Appendix illustrates the difference in goal diversity among datasets.
@@ IV-E Robot Deployment
We deploy LiMo in closed loop on an ANYmal D quadruped robot in locations outside the training scenes. These experiments aim to validate usability in real-world closed-loop control (Q4), embodiment-specific behavior (Q3), and robustness in unseen, cluttered environments (Q2). We showcase LiMo clearing obstacles, following corridors, and quickly reacting to dynamic scene changes.
We do not include on-robot comparisons to prior end-to-end visual navigation models. Existing approaches are designed for cross-embodiment deployment and often do not support pose-based goal conditioning. In contrast, LiMo predicts high-resolution, goal-pose-conditioned $\mathrm{SE}(2)$ trajectories that terminate at the commanded goal and are tailored to ANYmal’s embodiment. Moreover, representative models such as OmniVLA [15] are substantially larger, cannot be executed at comparable rates on the robot’s onboard hardware, and rely on additional mechanisms such as topological maps to provide subgoals. LiMo operates without such infrastructure.
Due to these fundamental differences in design objectives and deployment assumptions, real-robot comparisons are not directly interpretable and are therefore omitted.
@@ IV-E1 Deployment Setup
We run LiMo on a NVIDIA Jetson Orin on-board the robot in closed loop at $6\text{\,}\mathrm{H}\mathrm{z}$. We use a simple lookahead path follower node to make the robot track the predicted $\mathrm{SE}(2)$ waypoints. No geometric planning or sensing is used during the deployment. LiDAR point clouds are recorded solely to improve visualizations.
@@ IV-E2 Incorporating Side Cameras
Using only the front-facing camera in closed-loop deployment led to collisions, as obstacles were no longer detected once they left the forward field of view. Therefore, in the robot experiments, we use a multi-camera variant with additional left- and right-facing cameras, implemented via a minimal architectural extension that does not introduce memory or geometric sensing. A more detailed discussion of this limitation is given in Sec. -C.
@@ IV-F Obstacle Course Navigation
Figure 6: Robot navigating an indoor obstacle course at four different time steps. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory is shown in cyan and the goal is shown in green. LiMo generates smooth, collision-free paths while adapting to obstacles.
@@ IV-G Reacting to Dynamic Obstacles
Figure 7: The robot reacts to a dynamic obstacle (a person carrying a red mat). Rows show successive time steps during execution. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory (cyan) adapts online as the obstacle blocks the original path, leading to safe avoidance behavior.
To test reactivity, we introduce a moving obstacle by having a person carry a large mat and intentionally block the robot’s path during execution. The planner runs continuously at $6\text{\,}\mathrm{H}\mathrm{z}$, allowing the robot to replan as the obstacle moves.
@@ V Limitations
Planner supervision depends on elevation and traversability estimates and can fail in edge cases, leading to unsafe or suboptimal labels. At inference, LiMo operates on a single RGB frame without temporal memory, which limits reasoning under occlusions and causes obstacles to be forgotten once they leave the field of view. While geometric augmentation scales supervision efficiently, visual diversity is still constrained by the underlying dataset. Fig. 4 illustrates representative failure cases, and a continued discussion is provided in Sec. -D.
@@ VI Conclusion
We presented LiMo, a goal-conditioned visual navigation policy trained on a mix of teleoperation data and scalable geometric supervision. By combining expert demonstrations with planner-generated trajectories that are fully embodiment-specific, we scale supervision to thousands of hours of high-quality trajectories. Across held-out missions, geometric augmentation leads to clear improvements in SPL, especially in diverse, obstacle-rich environments. LiMo learns geometry-aware, semantically grounded behaviour aligned with ANYmal’s embodiment, and experiments on a real robot validate its utility as a practical path planner.
Overall, our results show that for vision-based goal-conditioned navigation, curating diverse, geometry-aware supervision can be more effective than relying solely on large, heterogeneous datasets. By combining limited expert demonstrations with scalable, embodiment-specific geometric supervision, we provide a practical and data-efficient path toward robust visual navigation in real-world robotic systems.
@@ Acknowledgments
Fig. 8. Visualization of goal distributions based on 10,000 samples from each dataset. (Left) Goals from the teleoperation dataset $\mathcal{D}_{\text{TEL}}$ (blue), sampled along the teleoperated trajectories. (Right) Goals from the augmented dataset $\mathcal{D}_{\text{AUG}}$, including teleoperation goals (blue) and additional randomly sampled goals (orange).
@@
Figure 10: The robot is following a narrow corridor toward a goal located at the far end. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory (cyan) remains centered and collision-free despite tight lateral constraints.

### v2

@@ IV-D Ablations
Figure 5: (Left) Evaluation of SPL on $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{AUG}}$ test sets trained on a varying number of augmenting geometric paths. Adding geometric samples mainly benefits performance on the more diverse $\mathcal{D}_{\text{AUG}}$ test split. (Middle) When evaluating goals following the $\mathcal{D}_{\text{AUG}}$ distribution, policies trained solely on $\mathcal{D}_{\text{TEL}}$ consistently underperform compared to those trained on the $\mathcal{D}_{\text{AUG}}$. While training on $\mathcal{D}_{\text{AUG}}$ generalizes well to $\mathcal{D}_{\text{TEL}}$-style goals, policies trained only on $\mathcal{D}_{\text{TEL}}$ fail to generalize to unseen diverse goals with limited data. (Right) Evaluation of SPL on $\mathcal{D}_{\text{TEL}}$ goals trained using the $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{AUG}}$. Training on just 5 missions of $\mathcal{D}_{\text{AUG}}$ achieves comparable performance to training on $\mathcal{D}_{\text{TEL}}$ with 41 missions.
@@ IV-D1 Training with fewer Geometric Paths per Frame
To explore the benefits of adding geometric paths, we train policies on different mixtures of $\mathcal{D}_{\text{TEL}}$ and $\mathcal{D}_{\text{GEO}}$. The policies are evaluated on the held-out test missions with $\mathcal{D}_{\text{TEL}}$ or more diverse $\mathcal{D}_{\text{AUG}}$ distributed goals (see Fig. 5 (left)).
@@ IV-D2 Training on fewer Missions
To understand how well policies generalize when trained with fewer missions, we compare performance across varying numbers of training missions drawn from either $\mathcal{D}_{\text{TEL}}$ or $\mathcal{D}_{\text{AUG}}$. Policies trained on a small number of diverse $\mathcal{D}_{\text{AUG}}$ missions retain strong performance on $\mathcal{D}_{\text{TEL}}$ goals, whereas policies trained only on $\mathcal{D}_{\text{TEL}}$ require substantially more missions to reach comparable SPL (see Fig. 5). This highlights that diversity in the training data, rather than sheer quantity, is key for efficient generalization (Q1, Q2). Tab. III in the Appendix illustrates the difference in goal diversity among datasets.
@@ IV-E Robot Deployment
We deploy LiMo on an ANYmal D quadruped robot in locations outside the training scenes to demonstrate its usability in real-world closed-loop control (Q4), embodiment-specific behavior (Q3), and robustness in unseen, cluttered environments (Q2). These experiments illustrate the robot clearing obstacles, following corridors, and reacting to dynamic scene changes.
We omit on-robot comparisons to prior end-to-end visual navigation models. Existing approaches are designed for cross-embodiment deployment and often do not support pose-based goal conditioning. In contrast, LiMo predicts high-resolution, goal-pose-conditioned $\mathrm{SE}(2)$ trajectories that terminate at the commanded goal and are tailored to ANYmal’s embodiment.
Moreover, representative models such as OmniVLA [12] are substantially larger, cannot be executed at comparable rates on the robot’s onboard hardware, and rely on additional mechanisms such as topological maps to provide subgoals. We therefore deliberately choose to only highlight the real-world performance of our most powerful model.
@@ IV-E1 Deployment Setup
We run LiMo on a NVIDIA Jetson Orin on-board the robot in closed loop at $6\text{\,}\mathrm{H}\mathrm{z}$. We use a simple lookahead path follower node to make the robot track the predicted $\mathrm{SE}(2)$ waypoints.
@@ IV-E2 Incorporating Side Cameras
Using only the front-facing camera in closed-loop deployment led to collisions, as obstacles were no longer detected once they left the forward field of view. Therefore, in the robot experiments, we use a multi-camera variant with additional left- and right-facing cameras, implemented via a minimal architectural extension.
More details are provided in Sec. -C.
@@ IV-F Obstacle Course Navigation
Figure 6: LiMo navigating an indoor obstacle course shown at four different time steps. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory is shown in cyan and the goal is shown in green. LiMo generates smooth, collision-free paths while adapting to obstacles. The LiDAR data is recorded for visualization purposes only and is not used as input to the policy.
@@ IV-G Reacting to Dynamic Obstacles
Figure 7: LiMo reacting to a dynamic obstacle (a person carrying a red mat). Rows show successive time steps during execution. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory (cyan) adapts online as the obstacle blocks the original path, leading to safe avoidance behavior.
To test reactivity, we introduce a moving obstacle by having a person carry a large mat and intentionally block the robot’s path during execution. LiMo runs continuously at $6\text{\,}\mathrm{H}\mathrm{z}$, allowing the robot to replan as the obstacle moves.
@@ V Limitations
Planner supervision depends on elevation and traversability estimates and can fail in edge cases, leading to unsafe or suboptimal labels. At inference, LiMo operates on a single RGB frame without temporal memory, which limits reasoning under occlusions and causes obstacles to be forgotten once they leave the field of view. While geometric augmentation scales supervision efficiently, visual diversity is still constrained by the underlying dataset. Fig. 4 illustrates representative failure cases, and a continuing discussion is provided in Sec. -D.
@@ VI Conclusion
We presented LiMo, a goal-conditioned visual navigation policy trained on a mix of teleoperation data and scalable geometric supervision. By combining expert demonstrations with planner-generated trajectories that are fully embodiment-specific, we scale supervision to thousands of hours of high-quality trajectories. Across held-out missions, geometric augmentation leads to clear improvements in SPL, especially in diverse, obstacle-rich environments. LiMo learns geometry-aware, semantically grounded behavior aligned with ANYmal’s embodiment, and experiments on a real robot validate its utility as a practical path planner.
Overall, our results show that for vision-based goal-conditioned navigation, curating diverse, geometry-aware supervision can be more effective than relying solely on large, heterogeneous datasets.
By combining limited expert demonstrations with scalable, embodiment-specific geometric supervision, we provide a practical and data-efficient path toward robust visual navigation in real-world robotic systems.
@@ Acknowledgments
Fig. 8. Visualization of goal distributions based on 10,000 samples from each dataset. (Left) Goals from the teleoperation dataset $\mathcal{D}_{\text{TEL}}$ (blue), sampled along the recorded trajectories. (Right) Goals from the augmented dataset $\mathcal{D}_{\text{AUG}}$, including teleoperation goals (blue) and additional randomly sampled goals (orange).
@@
Figure 10: LiMo is following a narrow corridor toward a goal located at the far end. From left to right: (i) third-person view, (ii) bird’s-eye view of the LiDAR point cloud with predicted paths, and (iii) left-, front-, and right-facing camera views. The predicted trajectory (cyan) remains centered and collision-free despite tight lateral constraints.